How many milliliters of pure acid must be added to 150 milliliters of a 70% solution of acid to obtain a 80% solution?

Answer by robertb(5830)

150*0.70 + x*1.00 = (150+x)*0.80
==> 105 + x = 120 + 0.80x
==> x - 0.80x = 120 - 105
==> 0.20x = 15
==> x = 75 milliliters of pure acid must be added.